# Scalable Multi-Agent Offline Reinforcement Learning and the Role of Information

### v1

@@ 3 Scalable Offline MARL
The proof of Lemma 3.2 builds on the simplifying assumption that at every step $k$ it is possible to (union) bound the error induced by the optimal solution $\|\tilde{q}^{k}-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\|^{2}_{2,\nu}\leq\epsilon_{\text{INH}}$ with $\tilde{q}^{k}=\operatorname*{arg\,inf}_{f\in\mathcal{F}_{i}}\|f-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\|^{2}_{2,\nu}$ and $\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}$ being the Bellman optimum. Interestingly, it is again possible to leverage the structure induced by the information-sharing network $\mathcal{G}$ to recover the following:
Lemma 3.4 (Th.1 of [3]).
Consider a fixed functional space $\mathcal{F}_{i}$ supported on $\mathcal{S}_{\mathcal{N}_{i}}$ induced by $\mathcal{G}$ and a full representation $\mathcal{F}_{i}^{\star}$ supported on $\mathcal{S}$, then by defining the Conditional Mutual Information (CMI) between two representations $\mathbb{I}(\mathcal{N}_{i})=\mathbb{I}(Y;s_{\mathcal{N}_{-i}}|s_{\mathcal{N}_{i}})$, with $Y$ being the target of the regression problem, and by defining the irreducible error under the full representation $\sigma^{2}=\left\|\tilde{q}^{k,\star}-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\right\|_{2,\nu}^{2}$, it follows that $\|\tilde{q}^{k}-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\|^{2}_{2,\nu}\leq\sigma^{2}+\mathbb{I}(\mathcal{N}_{i}).$
This Lemma states that by employing reduced-size information, one can bound the resulting error by the CMI between the target and the unused states, conditioned on the used ones. In other words, whenever a subset of states has low relevancy or high redundancy with the ones employed, that is, a low CMI, such states can be safely removed from the information-sharing graph between agents without significantly affecting the performance guarantees of the SCAM-FQI (or any other algorithm based on the same kind of supervised-learning step).
@@ 4 Empirical Corroboration: Multi-Agent Production Scheduling
Figure 1: $4\times 5$ Densely Connected Agents Layout. The time (in ticks) required to perform all the required operations is plotted against SCAM-FQI iterations (lower is better). The legend provides insights on the values of $d$ followed by whether the shared information was full (f) or compressed (c). Confidence intervals are bootstrapped $95\%$ confidence interval over 5 seeds.
@@
As expected from the analysis in the previous Section, both the reported experiments show a relevant and common trend: the information-sharing among agents does induce a trade-off between the quality of the learned policies and the speed of convergence; a higher degree of sharing improves the performances while compressed forms of sharing allows for faster convergence while not harming the overall performances; however, over-sharing ends up causing a degradation of performances.
@@ 5 Conclusions
In this work, we first introduced a novel routine for multi-agent offline data collection, by building different per-agent datasets based on an information-sharing graph rather than centralized or decentralized datasets. We then proposed an algorithm, SCAM-FQI, that takes explicit advantage of such structure to learn scalable decentralized policies. Furthermore, we provided theoretical guarantees of SCAM-FQI convergence in terms of the overall quality of the information shared between the agents. Finally, we provided some empirical corroboration of the trade-off induced by the quality of the information sharing by testing SCAM-FQI against a distributed production scheduling problem.

### v2

@@ 3 Towards Scalable Offline MARL
The proof of Lemma 3.2 builds on the simplifying assumption that at every step $k$ it is possible to (union) bound the error induced by the optimal solution $\|\tilde{q}^{k}-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\|^{2}_{2,\nu}\leq\epsilon_{\text{INH}}$ with $\tilde{q}^{k}=\operatorname*{arg\,inf}_{f\in\mathcal{F}_{i}}\|f-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\|^{2}_{2,\nu}$ and $\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}$ being the Bellman optimum. Yet, it is again possible to leverage the structure induced by the information-sharing network $\mathcal{G}$ to recover the following:
Lemma 3.4 (Th.1 of Beraha2019FeatureSV ).
Consider a fixed functional space $\mathcal{F}_{i}$ supported on $\mathcal{S}_{\mathcal{N}_{i}}$ induced by $\mathcal{G}$ and a full representation $\mathcal{F}_{i}^{\star}$ supported on $\mathcal{S}$, then by defining the CMI between two representations $\mathbb{I}(\mathcal{N}_{i})=\mathbb{I}(Y;s_{\mathcal{N}_{-i}}|s_{\mathcal{N}_{i}})$, with $Y$ being the target of the regression problem, and by defining the irreducible error under the full representation $\sigma^{2}=\left\|\tilde{q}^{k,\star}-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\right\|_{2,\nu}^{2}$, it follows that $\|\tilde{q}^{k}-\bar{\mathcal{T}}^{i}\hat{q}^{k-1}_{i}\|^{2}_{2,\nu}\leq\sigma^{2}+\mathbb{I}(\mathcal{N}_{i}).$
This Lemma states that by employing reduced-size information, one can bound the resulting error by the CMI between the target and the unused states, conditioned on the used ones. In other words, whenever a subset of states has low relevancy or high redundancy with the ones employed, that is, a low CMI, such states can be safely removed from the information-sharing graph without significantly affecting the performance guarantees of the SCAM-FQI.
@@ 4 Empirical Corroboration: Multi-Agent Production Scheduling
Figure 1: $4\times 5$ Densely Connected Agents Layout. The time (in ticks) required to perform all the required operations is plotted against SCAM-FQI iterations (lower is better). The legend provides insights on the values of $d$ followed by whether the shared information was full (f) or compressed (c). Confidence intervals are bootstrapped $95\%$ CI over 5 seeds.
@@
As expected from the analysis in the previous Section, both the reported experiments show a relevant and common trend: the information-sharing among agents does induce a trade-off between the quality of the learned policies and the speed of convergence; a higher degree of sharing improves the performances while compressed forms of sharing allows for faster convergence while not harming the overall performances; however, over-sharing ends up causing a degradation of performances: the shared information increases the variance of the dataset without adding any value in terms of reduced bias, eventually affecting the learning negatively.
@@ 5 Conclusions
In this work, we first introduced a novel routine for multi-agent offline data collection, by building different per-agent datasets based on an information-sharing graph rather than centralized or decentralized datasets. We then proposed an algorithm, SCAM-FQI, that takes explicit advantage of such structure to learn scalable decentralized policies. Furthermore, we provided theoretical guarantees of SCAM-FQI convergence in terms of the overall quality of the information shared between the agents. Finally, we provided some empirical corroboration of the trade-off induced by the quality of the information sharing by testing SCAM-FQI against a distributed production scheduling problem. We believe that our work can be a crucial step in the direction of developing scalable solutions in a principled way to yet more practical settings.
Comments on Practical Applicability.
The proposed approach integrates a decentralized architecture with a structured communication network. This design mitigates the need for full-state observability at deployment, offering a clear advantage over fully centralized solutions. However, it presupposes the availability of an efficient communication protocol in fully decentralized environments. Consequently, the practical applicability of the method is best suited to high-dimensional, communication-efficient scenarios; settings where centralized approaches struggle to scale, but where a communication infrastructure can be feasibly deployed.